### v2

# Is Large-scale Pretraining the Secret to Good Domain Generalization?

Abstract
Multi-Source Domain Generalization (DG) is the task of training on multiple source domains and achieving high classification performance on unseen target domains. Recent methods combine robust features from web-scale pretrained backbones with new features learned from source data, and this has dramatically improved benchmark results. However, it remains unclear if DG finetuning methods are becoming better over time, or if improved benchmark performance is simply an artifact of stronger pre-training. Prior studies have shown that perceptual similarity to pre-training data correlates with zero-shot performance, but we find the effect limited in the DG setting. Instead, we posit that having perceptually similar data in pretraining is not enough; and that it is how well these data were learned that determines performance. This leads us to introduce the Alignment Hypothesis, which states that the final DG performance will be high if and only if alignment of image and class label text embeddings is high. Our experiments confirm the Alignment Hypothesis is true, and we use it as an analysis tool of existing DG methods evaluated on DomainBed datasets by splitting evaluation data into In-pretraining (IP) and Out-of-pretraining (OOP). We show that all evaluated DG methods struggle on DomainBed-OOP, while recent methods excel on DomainBed-IP. Put together, our findings highlight the need for DG methods which can generalize beyond pretraining alignment.
1 Introduction
Domain Generalization (DG) addresses the challenge of enabling AI models to generalize from known domains to unseen ones, a critical task given the inevitable distribution shifts between training and real-world deployment (Saenko et al., 2010). DG pipelines typically consist of three stages: pretraining a model on a large, general dataset; finetuning the model with one or more source domains; and finally evaluating the model on target domains that are distinct from source domains. More and more, DG methods rely on huge-scale foundation models for initialization. Simultaneously, finetuning has increasingly incorporated regularization to prevent catastrophic forgetting. As a result, it remains unclear whether DG adaptation methods are genuinely improving or if enhanced benchmark performance is simply due to stronger pre-training combined with regularization, or even the presence of target domains within the hundred million-scale pre-training data.
In this work, we examine the reliance of recent DG methods on pre-trained features. We focus on CLIP-based models, as they are used in all state-of-the-art DG methods (Addepalli et al., 2024; Cho et al., 2023; Shu et al., 2023; Mao et al., 2024). While prior studies  (Mayilvahanan et al., 2024) have shown that perceptual similarity to pre-training data explains zero-shot performance—referred to as the Image Similarity Hypothesis—we find this relationship to be limited in the DG setting. Despite evidence of target domains being present in pre-training (Figure 6), perceptual similarity alone does not fully explain accuracy in the DG context. We propose that it is not just the presence of similar data in pre-training that matters, but also how well this data was learned. To this end, we introduce the Alignment Hypothesis, which states that pre-trained alignment between image and class embeddings is still predictive of DG performance even after source finetuning. We find that, performance for low alignment samples can be almost 0, while performance for high alignment samples is close to perfect. These results confirm the Alignment Hypothesis. As illustrated in Figure 1, these findings suggest that current DG methods largely fail to learn new, general features from the source data when the pretraining does not already provide a strong alignment.
Figure 1: An overview of desired and actual behaviour of DG methods. 1) DG methods are initialized with foundation models like CLIP. Pre-trained embeddings are relatively well aligned with ground truth labels on both source and target data for most samples (In-Pretraining, IP), but some samples are not well aligned (Out-of-pretraining, OOP). 2) An ideal DG method would strengthen alignment for both OOP and IP data with ground truth labels. 3) Our analysis shows that DG methods only result in strong alignment for IP data, leaving OOP data misaligned (Figure 2).
The confirmation of the Alignment Hypothesis gives us a tool to separate in-pretraining (IP) data from out-of-pretraining (OOP) data for a particular backbone, and we do so for five DG datasets with OpenCLIP-ViT/B-16. We call the resulting splits DomainBed-IP and DomainBed-OOP. Evaluating on DomainBed-IP/OOP offers a view of where current DG methods fail and where they succeed. We find that all methods, including those considered state-of-the-art, perform poorly on OOP data, i.e. data that the pretrained backbone hadn’t already aligned well. Furthermore, recent state-of-the-art methods do not outperform older methods on OOP data. For example, CLIPood  (Shu et al., 2023) slightly under performs a combination of older methods (MIRO (Cha et al., 2022) + MPA (Arpit et al., 2022)) on DomainBed-OOP. At the same time, existing DG methods show exceptional performance on DomainBed-IP, sometimes even outperforming an oracle model trained on the target domain. These results suggest that future research should aim to enhance DG methods on low-alignment data while preserving the already strong performance on high-alignment data. In summary, we make the following contributions:
  • Introduce the Alignment Hypothesis: We demonstrate that pre-training alignment between image and class text embeddings is a stronger predictor of Domain Generalization (DG) success than the previously proposed Image Similarity Hypothesis  (Mayilvahanan et al., 2024). Based on this, we define In-Pretraining (IP) as data well-aligned with pre-trained embeddings, and Out-of-Pretraining (OOP) as data with weaker alignment.
  • Propose a new IP/OOP evaluation framework: We demonstrate that splitting target data by its alignment with the pre-trained backbone can effectively test Out-of-Pretraining (OOP) generalization. We will release IP/OOP splits for the DomainBed dataset to support future research.
  • Expose strengths and limitations of state-of-the-art DG methods: Using DomainBed-IP/OOP we find that leading DG methods perform well on data well-aligned by pre-training but struggle on misaligned samples, emphasizing the need for methods that move beyond reliance on pre-training.
2 Related Work
Multi-Source Domain Generalization: Domain Generalization aims to mitigate the impacts of domain shifts between source (training) and target (deployment) domains. One standard approach is domain-invariant feature learning, which leverages domain labels to learn domain-invariant features. CORAL  (Sun & Saenko, 2016) aligns second-order statistics, while DANN  (Ganin et al., 2016) uses an adversarial loss.  Gulrajani & Lopez-Paz (2020) show that ERM, which does not align features between domains, can outperform most prior work while being easier to tune. Another common approach is domain-aware data augmentation to expand the training domain to become closer to or even overlap the target domain. Inter-domain mixup  (Yan et al., 2020) blends images from different domains. Similarly, style transfer can diversify training images  (Zhong et al., 2022). Deep ensembles are effective for domain generalization (Arpit et al., 2022). Since they are computationally inefficient for inference, many recent works average model weights from either multiple finetuning runs or from a single training trajectory (Cha et al., 2021; Arpit et al., 2022; Rame et al., 2022; Jain et al., 2023; Li et al., 2023; Shu et al., 2023). More recently, several methods perform regularized finetuning towards the initialization of a pretrained model. This works under the assumption that pretrained features are useful for target data, and should not be unlearned. The general idea can be applied to weight space (L2SP (Xuhong et al., 2018)), feature space (MIRO (Cha et al., 2022)), or output space (CAR-FT (Mao et al., 2024),CLIPood (Shu et al., 2023)).
Large-Scale Pretraining for DG: Recent DG literature (Cho et al., 2023; Cha et al., 2022; Addepalli et al., 2024; Mao et al., 2024; Arpit et al., 2022) leverages large-scale pretrained initializations stronger than ImageNet (Russakovsky et al., 2015), and CLIP (Radford et al., 2021) is the most common choice. CLIP leverages a cross-domain contrastive loss to align images and captions. Due to the large scale of training data (typically at least 400 million samples) and the free-form nature of the text, CLIP enables effective zero-shot classification and learns features that generalize very well. Other choices for very strong pretraining include SWAG (Singh et al., 2022) and DinoV2 (Oquab et al., 2023). SWAG uses supervision from Instagram hashtags, while DinoV2 is trained without text supervision and instead relies on augmentation-based alignment. While our analysis focuses on image-text models like CLIP due to its popularity, the concept of alignment can extend to other types of pretraining models. We leave the exploration of this extension to future work.
Impact of Data on Model Performance: Several recent studies have explored the influence of pretraining data on model performance. Mayilvahanan et al. (2024) investigated how the presence of perceptually similar images in CLIP (Radford et al., 2021) pretraining affects performance, introducing the Similarity Hypothesis, which posits that nearest neighbor similarity is strongly correlated with zero-shot accuracy. Udandarao et al. (2024) demonstrated that concept frequency in pretraining is correlated with zero-shot performance and introduced a dataset focusing on infrequent concepts. Fang et al. (2022) found that diversity in pretraining data is critical for improving performance on benchmarks such as ImageNetV2 (Recht et al., 2019), ImageNet-R (Hendrycks et al., 2021), ImageNet-Sketch (Wang et al., 2019), and ObjectNet (Barbu et al., 2019). However, these studies focus on the zero-shot setting, where models are evaluated without further training. In contrast, we examine the domain generalization setting, where pre-trained models are fine-tuned on source domains and tested on held-out target domains. Our findings suggest that comparing target images to pre-trained images, as proposed by Mayilvahanan et al. (2024), is less predictive of final DG performance than directly measuring the alignment between the image and its class embedding.
3 Analyzing the role of pretraining in Domain Generalization
This work explores Multi-Source Domain Generalization for classification, where samples from multiple source domains (e.g., sketches, product photos) and a held-out target domain (e.g., wildlife camera images) are annotated with both domain and class labels. We construct a training dataset by aggregating all sample-label pairs from all training domains $d\in\{d_{1},\ldots,d_{n}\}$, denoted as
$D=\{(X^{d_{1}},Y^{d_{1}}),\ldots,(X^{d_{n}},Y^{d_{n}})\}.$
We initialize a classifier $f$ with a contrastively pre-trained vision-language model (e.g., CLIP) and finetune it on $D$. The scale of pre-training datasets is many orders of magnitude larger than that of source datasets. Most methods fully fine-tune $f$, though LP-FT (Kumar et al., 2021) fine-tunes the linear probe before the main network and Attention Tuning (Teterwak et al., 2023; Touvron et al., 2022) only tunes attention layers. The performance is then evaluated on a held-out testing domain $d_{\text{test}}$. The key assumption is that $d_{test}$ has a different distribution from the source domains.
We aim to analyze how reliant existing DG methods are on pre-training. A recent analysis of CLIP proposed the Image Similarity Hypothesis  (Mayilvahanan et al., 2024), which supposes that high CLIP performance on a given test sample is a result of highly similar nearest-neighbor images in pre-training, and tested it on zero-shot classification tasks. They found a strong correlation between nearest-neighbor similarity and zero-shot classification performance, but did not analyze OOD performance after fine-tuning. Therefore, we apply an equivalent testing setup for the DG setting, where a pre-trained model is fine-tuned on a source distribution and tested on a different target distribution. We find only a limited influence of image similarity in Section 3.1. To better understand the role of pretraining in domain generalization, we introduce the Alignment Hypothesis, which we explore in detail in Section 3.2. We later use the Alignment Hypothesis to split DG datasets and analyze existing DG methods (Section 4).
3.1 Testing the Image Similarity Hypothesis
The Image Similarity Hypothesis  (Mayilvahanan et al., 2024) posits that test performance improves when there are perceptually similar images in the pre-training dataset. The PerceptualSimilarityScore measures perceptual similarity and is defined as the cosine similarity between a target image $I$ and its nearest neighbor $I_{k}$ in pre-training:
[…]
Figure 2: Comparing the Predictive Power of the Alignment and Image Similarity Hypotheses for Domain Generalization (DG). a.) Alignment Hypothesis: The cosine similarity between image and ground truth text-label embeddings before pre-training (Alignment Score) is highly predictive of model accuracy after fine-tuning on five DG datasets, with Alignment Score distributions shown in the colored histograms. This suggests that image-text pairs well-aligned during pre-training result in better performance on target tasks. b.) Image Similarity Hypothesis: In contrast, the cosine similarity between a test image and its closest match from the pre-training set (Perceptual Similarity Score) shows weaker predictive power for accuracy, implying that visual resemblance alone is less indicative of downstream performance.
[…]
3.2 Introducing the Alignment Hypothesis
To find a stronger predictor of DG accuracy than perceptual similarity, we focus on how effectively pre-training captures the relationship between an image and its label. This leads us to propose the Alignment Hypothesis, which states that if an input image and its corresponding text label (e.g., ‘A photo of a {cls}’) are well-aligned in the embedding space, final DG performance will be high. Crucially, alignment is measured before source fine-tuning while DG performance is measured after adaptation. This allows us to isolate the contribution of fine-tuning. Since models like CLIP optimize image-text pairs using a contrastive loss, cosine similarity between image and text embeddings is an alignment measure well coupled to their training objective. Therefore, we use it as our metric of pre-training generalization. More formally:
$\text{AlignmentScore}(I,T)=\frac{\langle f_{I}(I),f_{T}(T)\rangle}{\|f_{I}(I)\|\cdot\|f_{T}(T)\|}$
where $f_{I}(I)$ is the embedding of the image before finetuning on source, and $f_{T}(T)$ is the embedding of the text.
We verify the Alignment Hypothesis similarly to the Image Similarity Hypothesis, by binning samples using the AlignmentScore and computing accuracy for each bin using VL2V-SD. We provide the same analysis for many more DG methods in the Appendix Figure 8. In Figure 2 a.), we can see that the Alignment Hypothesis explains DG performance after source finetuning, significantly more strongly than for the Image Similarity Hypothesis in Figure 2 b.) This finding suggests that source fine-tuning in DG, which aims to achieve high performance across all target samples, only succeeds on those with high initial alignment.
4 Re-thinking Domain Generalization Benchmarking using the Alignment Hypothesis
Knowing that the Alignment Hypothesis holds for contrastively trained image-text models (Section 3), we can now use it as a tool to probe the performance of DG methods across different levels of pre-training alignment. We apply this approach to five widely-used DomainBed (Gulrajani & Lopez-Paz, 2020) DG datasets: VLCS (Fang et al., 2013), PACS (Li et al., 2017), OfficeHome (Ganin et al., 2016), TerraIncognita (Beery et al., 2018), and DomainNet  (Peng et al., 2019). This section discusses how we create new splits for existing DG datasets using AlignmentScore. We start by computing AlignmentScore for all samples in 5 DG datasets (Figure 3). Based on our observation that some samples are mislabeled, we perform dataset cleaning (4.1). Finally, we find an AlignmentScore threshold to split DG datasets into In-Pretraining (IP) and Out-of-Pretraining (OOP) evaluation subsets (Section 4.2), which we later use for evaluating DG methods (Section 5). In order to connect the AlignmentScore with the DG method, we use the same backbone both for splitting the datasets into IP and OOP subsets and for training DG methods.
Figure 3: Representative DomainBed dataset samples and their labels at various AlignmentScore values. At very low AlignmentScores, most labels (red boxes) are incorrect. At very high AlignmentScores, text present in the image corresponds to the label.
4.1 Data Exploration and Cleaning
We start by visualizing the data of all the datasets at various AlignmentScore values. We show some representative samples in Figure 3. At very low scores, we find that a large fraction of labels are incorrect (red boxes in Figure 3).
As a result, we divide the data into AlignmentScore intervals (e.g., 0.00-0.05, 0.05-0.10, and so on, up to 0.2) and randomly sample 100 instances from each interval for every dataset. This allows us to systematically analyze the relationship between AlignmentScore and label accuracy across different score ranges. For each interval, we then count the fraction of mislabelled samples to better understand how low AlignmentScores are associated with labeling errors. We find that below an AlignmentScore of 0.15, label noise is unacceptably high, with all datasets suffering the most from mislabelling (Table 4 in Appendix). Therefore, we discard all samples with AlignmentScore less than 0.15 in DomainBed-IP/OOP. As shown in Table 1, we observe that the percentage of discarded samples due to mislabeling varies across datasets, with VLCS and DomainNet having the highest rates at 12.41% and 7.64%, respectively.
Furthermore, on the right side of Figure 3, we observe that at very high AlignmentScores (greater than 0.4), the images often contain text directly related to the label. Since our goal is to evaluate visual recognition rather than text recognition (OCR), and CLIP is known to have strong OCR abilities  (Fort, 2021), we exclude all samples above AlignmentScore of 0.4 from DomainBed-IP/OOP. As shown in Table 1, although only a small portion of data is removed due to OCR filtering (0.00-0.15% across datasets), this issue may become more significant in future studies.
Table 1: Percentage of discarded samples due to mislabeling or label text in the image (OCR).
Dataset Dropped - Mislabelled Dropped - OCR
OfficeHome 0.92% 0.15%
PACS 3.05% 0.00%
TerraIncognita 0.22% 0.00%
DomainNet 7.64% 0.03%
VLCS 12.41% 0.00%
4.2 Data Splitting
After filtering, we focus on determining a threshold to split the dataset into In-Pretraining (IP) and Out-of-Pretraining (OOP) subsets. We select 0.21 as the threshold, based on the trends observed in Figure 2 (a.)), as this is the point where performance begins to improve significantly, indicating that existing methods become more effective. While this threshold represents a somewhat subjective choice informed by observed patterns, we provide AlignmentScores in the released data, allowing researchers the flexibility to experiment with their own thresholds.
Figure 4 illustrates how this split impacts the size and composition of each dataset. Certain domains naturally fall into the IP subset. For example, in DomainNet, the clipart domain is predominantly categorized as IP, likely due to its frequent presence on the internet and therefore the web-scraped pre-training data. On the other hand, TerraIncognita-OOP has relatively more balanced domains, but exhibits substantial class shift between OP and OOP splits (Figure 5), meaning some classes are better aligned than others during pre-training. More dataset statistics are in the Appendix.
Figure 4: DomainBed-(IP/OOP) Statistics: Breakdown of DomainBed-IP and DomainBed-OOP counts, by dataset and domain. Overall, DomainNet and VLCS have the largest fraction of samples falling into DomainBed-OOP.
Figure 5: Class-distribution shift: TerraIncognita’s class distribution differs between DB-IP and DB-OOP, indicating that some classes were better aligned during pretraining.
5 Experiments
5.1 Training and Evaluation Protocol
We aim to evaluate how well Domain Generalization (DG) methods perform across both pretraining-aligned(IP) and pretraining-misaligned(OOP) data. We adhere to the DomainBed evaluation methodology, where one domain is chosen as the target, and the remaining domains act as source domains. To maintain a sufficient amount of training data, we train all DG methods on the original, unsplit datasets. We use hyper-parameter values recommended by the original implementation authors for each method.
After training, models are evaluated separately on the IP and OOP subsets, as well as the original, unsplit test domain. This approach allows us to measure how well each method generalizes to both pretraining-aligned and pretraining-misaligned data.
We follow the literature’s standard practice of computing performance per target domain and averaging the results across all domains. This method ensures that any domain imbalances do not disproportionately influence the final performance metrics.
Table 2: Benchmarking DG methods on DB-IP/OOP
DomainBed-IP DomainNet OfficeHome PACS TI VLCS Average
CORAL (Sun & Saenko, 2016) 63.3 76.1 84.3 42.9 86.5 70.6
SAGM (Wang et al., 2023) 64.3 79.5 90.1 44.0 88.0 73.2
ERM* (Gulrajani & Lopez-Paz, 2020) 63.1 78.1 87.1 42.0 85.3 71.1
LP-FT (Kumar et al., 2021) 64.4 78.5 90.3 40.9 86.0 72.0
SWAD (Cha et al., 2021) 72.3 84.8 94.6 52.7 88.5 78.6
MIRO (Cha et al., 2022) 72.4 88.8 97.6 58.9 91.0 81.7
VL2V-SD (Addepalli et al., 2024) 78.1 91.4 98.0 48.1 92.4 81.6
Attn Tune (Teterwak et al., 2023) 69.2 84.8 96.4 53.0 88.7 78.4
MPA (Arpit et al., 2022) 73.6 85.1 95.4 54.4 90.7 79.8
CLIPOOD (Shu et al., 2023) 78.9 90.9 97.7 63.5 92.5 84.7
MIRO + SWAD 77.0 90.5 97.6 62.1 91.1 83.6
MIRO + MPA 78.2 90.7 98.1 62.6 91.0 84.1
Upper Bound (Target Finetune) 81.6 88.5 97.8 93.4 93.8 91.0
(a) Samples with high AlignmentScore values, indicating good alignment between image and text features learned during pretraining.
DomainBed-OOP DomainNet OfficeHome PACS TI VLCS Average
CORAL (Sun & Saenko, 2016) 22.3 42.6 74.1 16.0 74.0 45.8
SAGM (Wang et al., 2023) 23.0 44.5 74.2 19.3 73.3 46.9
ERM* (Gulrajani & Lopez-Paz, 2020) 22.3 42.9 76.9 16.5 76.4 47.0
LP-FT (Kumar et al., 2021) 22.7 43.4 78.6 23.1 70.7 47.7
SWAD (Cha et al., 2021) 28.6 49.9 79.1 21.0 77.0 51.1
MIRO (Cha et al., 2022) 28.4 56.6 84.7 18.5 73.7 52.4
VL2V-SD (Addepalli et al., 2024) 31.8 56.6 85.0 15.9 79.1 53.7
Attn Tune (Teterwak et al., 2023) 26.8 51.4 84.2 20.3 76.1 51.8
MPA (Arpit et al., 2022) 29.6 51.0 82.7 22.2 79.5 53.0
CLIPOOD (Shu et al., 2023) 33.9 63.9 87.2 19.9 80.7 57.1
MIRO + SWAD 32.0 59.0 85.4 21.1 78.9 55.3
MIRO + MPA 33.1 60.0 87.8 24.9 80.3 57.2
Upper Bound (Target Finetune) 48.8 61.9 92.9 83.2 92.4 75.8
(b) Samples with lower AlignmentScore values, representing cases where pretraining alignment is weak.
DomainBed-All DomainNet OfficeHome PACS TI VLCS Average
CORAL (Sun & Saenko, 2016) 50.6 73.2 83.2 39.6 78.5 65.0
SAGM (Wang et al., 2019) 51.5 76.4 87.5 41.0 80.4 67.3
ERM* (Gulrajani & Lopez-Paz, 2020) 50.5 75.0 85.2 39.0 77.9 65.5
LP-FT (Kumar et al., 2021) 51.3 75.5 88.4 38.5 78.0 66.3
SWAD (Cha et al., 2021) 57.9 81.8 92.4 49.0 80.1 72.2
MIRO (Cha et al., 2022) 57.5 85.8 96.4 54.3 81.1 75.0
VL2V-SD (Addepalli et al., 2024) 62.0 88.3 96.9 44.4 82.7 74.9
Attn Tune (Teterwak et al., 2023) 55.4 81.9 95.4 49.1 81.8 72.7
MPA (Arpit et al., 2022) 58.9 82.0 94.3 50.7 82.3 73.6
CLIPOOD (Shu et al., 2023) 63.6 88.3 96.8 58.5 83.4 78.1
MIRO + SWAD 61.4 87.6 96.6 57.4 82.0 77.0
MIRO + MPA 62.4 87.9 97.2 58.2 82.8 77.7
Upper Bound (Target Finetune) 70.4 86.2 97.2 92.4 87.9 86.8
(c) Performance of DG methods on unsplit DomainBed
5.2 Algorithms
Explicit Regularization towards Pretraining: Several recent DG methods leverage explicit regularization towards the initialization. These methods generally operate either in weight space (by regularizing or freezing model parameters) or in feature space (by aligning internal feature representations with those of the pretrained model).
  • MIRO (Cha et al., 2022): Minimizes the Mutual Information between DG model intermediate features and CLIP intermediate features.
  • Attention Tuning (Teterwak et al., 2023; Touvron et al., 2022): Freezes all parameters except those in the Multiheaded-Attention Layers.
  • VL2V-SD (Addepalli et al., 2024): Self-distills a linear combination of CLIP vision and text outputs into a model.
  • CLIPood (Shu et al., 2023): Regularizes both weights and output features. Weights are averaged between the pre-trained CLIP model and the fine-tuned model, and outputs are regularized using a loss that incorporates class information from the pre-trained text encoder.
  • Linear Probe - Fine Tuning (LP-FT) (Kumar et al., 2021): Freezes the backbone, trains a linear probe, and then performs full finetuning. An untrained linear probe can cause finetuning to update the frozen backbone needlessly, potentially unlearning discriminative features. This biases the model towards pre-trained weights with smaller gradient updates.
Domain Invariance: A classic idea for Domain Generalization is Domain Invariance in the feature space, where the model learns only class-discriminative features shared among all training/source domains. CORAL (Sun & Saenko, 2016) matches the second moments of features across different domains and has been shown to be highly effective (Gulrajani & Lopez-Paz, 2020).
Flat Optima: Studies by Izmailov et al. (2018) and Cha et al. (2021) have shown that flat minima generalize better than sharp minima, as they make loss values less sensitive to perturbations in the loss surface, resulting in smaller increases in loss during domain shifts.
  • SWAD (Cha et al., 2021): Averages model parameters during training, determining the interval over which to average using validation loss over source domain.
  • Model Parameter Averaging (MPA) (Arpit et al., 2022): Starts averaging model parameters after a number of burn-in steps to find flat minima.
  • SAGM (Wang et al., 2023): An optimizer that explicitly optimizes for flat minima.
Baseline and Oracle: We also train a baseline and oracle model for lower-bound and upper-bound reference. The baseline model is an Empirical Risk Minimization (ERM) model that is finetuned on source domains and evaluated on target domains, and has been found to be effective for the DG task (Gulrajani & Lopez-Paz, 2020). The oracle model is trained on an 80% training split of all domains and evaluated on a 20% test split. The oracle model removes the OOD aspect of generalization and provides a reasonable upper bound for DG methods.
5.3 Results
We show results in Table 2(c) for DomainBed-OOP, DomainBed-IP, and standard DomainBed datasets. Underlined results represent the best performance of any single method (excluding method combinations), while the bold numbers show the highest performance overall, excluding the upper bound.
DG methods have nearly solved DomainBed-IP: In most datasets within DomainBed-IP, domain generalization approaches have nearly solved the problem of generalization. On three out of five datasets (DomainNet, OfficeHome, and PACS), the best DG method even outperforms the oracle! A notable exception is TerraIncognita, where CLIPood scores only 63.5%, far below the oracle’s 93.4%, highlighting this dataset as a challenging outlier. Interestingly, all three datasets where performance exceeds the oracle have an average IP AlignmentScore of 0.28, while the others have a lower average AlignmentScore of around 0.26. Therefore, the underperformance of TerraIncognita may be partially explained by its lower IP AlignmentScore, suggesting that alignment plays a significant role in DG performance, even within the IP case.
DG Methods leave much to be desired on DomainBed-OOP, but are still stronger than ERM: In DomainBed-OOP, we observe that even the top-performing DG methods struggle with low-alignment data. For example, CLIPood achieves 57.1% accuracy, which is a significant drop compared to its performance on DomainBed-IP (84.7%) and DomainBed-All (78.1%). Despite this, DG methods still outperform Empirical Risk Minimization (ERM), which scores only 47.0% on DomainBed-OOP. Therefore DG methods are better equipped to handle domain shifts than ERM, possibly due to weak transfer of knowlegdge from pre-trained features. Nevertheless, there is still substantial room for improvement on low-alignment samples.
State-of-the-Art methods do not consistently outperform older methods on OOP data: While CLIPood (Shu et al., 2023) clearly outperforms other methods on DomainBed-All with 78.1% accuracy, it performs comparably to older methods on DomainBed-OOP. For instance, MIRO + MPA (Cha et al., 2022; Arpit et al., 2022) achieves 57.2% on DomainBed-OOP, which is nearly the same as CLIPood’s 57.1%. This suggests that CLIPood’s primary advantage comes from well-aligned samples in DomainBed-IP, where it reaches 84.7% accuracy, half a percent better than the next-best method.
Model Parameter Averaging (MPA) boosts performance on OOP data: MPA performs significantly better than ERM on DomainBed-OOP, achieving 53.0% compared to ERM’s 47.0%. This is half a percent better than MIRO on OOP data, despite being 2% worse than MIRO on IP data. When combined with MIRO, MPA delivers the best performance on DomainBed-OOP, scoring 57.2%, slightly surpassing CLIPood. This suggests that MPA can complement other regularization-based methods like MIRO. On DomainBed-IP, MIRO + MPA scores 84.1%, close to CLIPood’s 84.7%, demonstrating versatility across both high- and low-alignment data. Interestingly, SWAD underperforms MPA on DomainBed-OOP by 2%, despite being conceptually similar. We attribute this to selecting the averaging interval on the source data, which introduces overfitting to source domains.
6 Conclusion
In this paper, we systematically explore how Domain Generalization (DG) methods rely on pre-trained feature alignment from models like CLIP. We hypothesize that the alignment between image and text embeddings during pre-training strongly predicts DG performance. Our experiments confirm this, showing that methods perform well on high-alignment samples (DomainBed-IP) but struggle on low-alignment data (DomainBed-OOP). While we focus on multi-source DG, we expect similar results in single-source finetuning due to the foundational role of pre-trained alignment. Notably, state-of-the-art methods like CLIPood perform near oracle-level on aligned data but see significant drops on misaligned samples. This suggests current DG methods rely on pre-trained features and fail to learn new, generalizable features from source domains. Moving forward, two paths emerge: developing DG methods that better learn generalizable features, or focusing on improving pre-trained backbones. While foundation models will continue to advance, there will always be specialized distributions where they fail. We hope these findings inspire further research into improving generalization on low-alignment data, pushing DG beyond reliance on pre-trained alignment.
DISTRIBUTION STATEMENT A. Approved for public release. Distribution is unlimited. This material is based upon work supported by the Under Secretary of Defense for Research and Engineering under Air Force Contract No. FA8702-15-D-0001. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Under Secretary of Defense for Research and Engineering.
Appendix A Appendix
A.1 Training and Evaluation Details
Code is zipped in the supplementary.
We use a slightly modified MIRO Cha et al. (2022) codebase for training and evaluation. We use leave-one-out evaluation, where a model is trained on all domains except the evaluation domain. We emphasize that we use DomainBed-IP and DomainBed-OOP as evaluation data only, models are trained on full datasets.
For training, we use an OpenCLIP-ViT-B/16 Ilharco et al. (2021) trained on LAION-400M Schuhmann et al. (2021). We use default hyper-parameters as defined by  Cha et al. (2022). This includes a learning rate of 5e-5, weight decay of 0.0, a batch size of 32 per-domain, an Adam Optimizer, and no dropout for all methods.
For evaluation, unlike DomainBed, we consider the entire test domain instead of an 80% random split. Following standard practice, we first compute accuracy for each domain, then average those accuracies to get dataset level statistics, and finally compute overall averages averaging across datasets.
For benchmarked methods, we also use hyper-parameters found to be best in respective papers. For SWAD, we use an optimum patience parameter value of 3, overfit patience parameter value of 6, and tolerance ratio of 6. For MIRO, we use use regularizer loss weight of 1.0. For CORAL, we use a CORAL regularizer weight of 1.0, following  Cha et al. (2021). For LP-FT, we train the linear probe for 600 steps before unlocking the full backbone. For Model Parameter Averaging, we burn in the training for 600 steps before averaging iterates. For VL2V-SD and CLIPood, we directly use the author’s implementation and hyper-parameters, except initializing with OpenCLIP Ilharco et al. (2021).
A.2 Training Compute
Each run uses an A6000 48GB GPU, trained for up to 12 hours per domain-dataset combination.
Appendix B Additional Results
B.1 AlignmentScore vs Accuracy
In Figure 8, we plot all benchmarked methods from the main paper, with x-axis corresponding to AlignmentScore, and the y-axis corresponding to the Top-1 Accuracy. We normalize for dataset size, so that no dataset dominates the count. In Figures 9 through 13, we plot these statistics independently per dataset, and find the trends consistent across datasets.
B.2 Per-Dataset Benchmarking Results
We expand Table 2(c) in the main paper into per-dataset results in Table 5 through 19.
B.3 Similarity of Target to Pre-training
To evaluate the Image Similarity Hypothesis, we retrieve the nearest neighbors from the Laion-400M dataset Schuhmann et al. (2021). This raises the question of how similar the target domains are to the pre-training data and whether the source domains might be even more similar. To investigate this, we compute Maximum Mean Discrepancy (MMD) distances between PACS domains and their nearest neighbors from Laion-400M, as shown in Figure 7. Our results indicate that target domains are, in fact, more similar to the pre-training data than source domains. We inspect nearest neighbors manually, and find even exact duplicates (Figure 6). Interestingly, while we found not only domain-level duplicates but also exact matches in the pre-training data, the Image Similarity Hypothesis is ultimately less predictive than the Alignment Hypothesis.
Figure 6: Nearest neighbors of target images in pre-training LAION data.
Figure 7: MMD between pre-training and target, source and target, and target and target for PACS. Target is more similar to pre-training than source. Despite this, Alignment is a better predictor of DG performance than perceptual similarity.
B.4 Other backbones
We benchmark 2 additional backbones(DINOv2 (Oquab et al., 2023) and OpenAI CLIP) using the MIRO + MPA Domain Generalization method, which we found to be the strongest in our paper, on two datasets (OfficeHome and PACS). This consistency is likely due to the similar nature of the pre-training datasets, both sourced from web scraping and of comparable scale. These findings reinforce that the usefulness of the DomainBed-IP/OOP split is not confined to a specific backbone.
Backbone OfficeHome - IP PACS - IP OfficeHome - OOP PACS - OOP
OpenCLIP-ViT/B 16 90.7 98.1 60.0 87.8
CLIP-ViT/B 16 88.1 97.6 57.0 86.9
DinoV2 87.4 97.1 58.9 85.0
Table 3: Benchmark results of different backbones on OfficeHome and PACS datasets.
Appendix C Additional DomainBed-(IP/OOP) Statistics and Analysis
C.1 Constituent Datasets
We split 5 datasets in DomainBed-(IP/OOP),( VLCS Fang et al. (2013), DomainNet Peng et al. (2019), OfficeHome Ganin et al. (2016), PACS Li et al. (2017), and TerraIncognita Beery et al. (2018)). Here we provide basic statistics of each.
VLCS has 5 classes and 4 domains: Caltech101, LabelME, SUN09, and VOC2007, with 10729 samples. The domain shift is dataset source.
DomainNet contrains 345 classes and 6 domains: clipart, infograph, quickdraw, real, and sketch. It has a total of 586,575 samples. The dataset shift is style.
OfficeHome has 65 classes and 4 domains: art, clipart, product, and real. The dataset shift is style.
TerraIncognita has 10 classes of wildlife cameras. There are 4 domains of different cameras and 24788 samples. The dataset shift is camera location.
PACS has 9991 sampes and 4 domains: arts, cartoon, photo, and sketch. There are 7 classes. The dataset shift is style.
C.2 Mislabeling Rates
In Table 4 we present the mislabelling rates at various AlingmentScore values.
Table 4: Mislabeling rates across different AlignmentScore ranges.
Dataset 0.0-0.05 0.05-0.10 0.10-0.15 0.15-0.20
OfficeHome 100% 45% 28% 12%
PACS - 46% 5% 0%
TerraIncognita - 55% 34% 23%
DomainNet 65% 35% 33% 11%
VLCS 33% 14% 9% 3%
C.3 Class Distribution of DomainBed-(IP/OOP):
In Figures 14 through 28, we provide class distribution statistics of different datasets before splitting and in our IP and OOP splits. We find some interesting patterns. For example, in Office-Home, the OOP class is dominated by marker and toys, while the IP split has a much more uniform distribution. Similarly, both PACS (DomainBed-OOP) and VLCS (DomainBed-OOP) are dominated by person.
Figure 8: Plotting AlignmentScore vs Top-1 Accuracy for all benchmarked methods, for all datasets together. Although some methods are stronger than others, all follow the same trend of increasing accuracy with increased AlignmentScore.
Figure 9: Plotting AlignmentScore vs Top-1 Accuracy on the PACS dataset.
Figure 10: Plotting AlignmentScore vs Top-1 Accuracy on the DomainNet dataset.
Figure 11: Plotting AlignmentScore vs Top-1 Accuracy on the OfficeHome dataset.
Figure 12: Plotting AlignmentScore vs Top-1 Accuracy on the TerraIncognita dataset.
Figure 13: Plotting AlignmentScore vs Top-1 Accuracy on the VLCS dataset.
Table 5: Per-domain breakdown for OfficeHome (DomainBed-OOP)
Method Art Clipart Product Real Avg
CORALSun & Saenko (2016) 33.0 29.4 50.0 58.0 42.6
SAGMWang et al. (2023) 36.8 30.7 51.7 58.7 44.5
ERM*Gulrajani & Lopez-Paz (2020) 36.0 28.1 51.3 56.0 42.9
LP-FTKumar et al. (2021) 37.2 29.7 52.0 54.7 43.4
SWADCha et al. (2021) 43.9 37.1 56.0 62.7 49.9
MIRO Cha et al. (2022) 49.4 39.6 62.7 74.7 56.6
VL2V-SD Addepalli et al. (2024) 56.4 39.3 65.3 65.3 56.6
Attn. Tune Teterwak et al. (2023); Touvron et al. (2022) 46.0 36.1 59.7 64.0 51.4
Model Parameter Averaging (MPA) Arpit et al. (2022) 46.0 36.4 55.7 66.0 51.0
CLIPOOD (Shu et al., 2023) 60.3 44.4 73.3 78.0 63.9
MIRO + SWAD 55.2 44.7 63.3 72.7 59.0
MIRO + MPA 53.5 44.1 65.7 76.7 60.0
Table 6: Per-domain breakdown for OfficeHome (DomainBed-IP)
Method Art Clipart Product Real Avg
CORAL 71.6 62.9 85.8 84.0 76.1
SAGM 75.7 70.8 87.0 84.5 79.5
ERM* 74.1 67.4 86.7 84.1 78.1
LP-FT 75.2 66.9 88.9 83.0 78.5
SWAD 81.4 77.8 91.8 88.2 84.8
MIRO 88.3 81.1 93.7 92.1 88.8
VL2V-SD 91.0 83.8 96.7 94.2 91.4
Attn. Tune 86.0 75.1 89.5 88.7 84.8
Model Parameter Averaging (MPA) 83.6 77.1 91.1 88.5 85.1
CLIPOOD (Shu et al., 2023) 92.3 81.3 96.0 94.2 90.9
MIRO + SWAD 90.4 83.2 95.5 92.8 90.5
MIRO + MPA 90.6 83.9 95.4 92.9 90.7
Table 7: Per-domain breakdown for OfficeHome (DomainBed-All)
Method Art Clipart Product Real Avg
CORAL 66.6 60.1 83.0 83.0 73.2
SAGM 70.7 67.1 84.3 83.6 76.4
ERM* 69.2 63.9 83.9 83.1 75.0
LP-FT 70.3 63.6 86.0 82.0 75.5
SWAD 76.6 74.3 89.0 87.2 81.8
MIRO 83.2 77.5 91.2 91.4 85.8
VL2V-SD 86.4 79.7 94.1 93.1 88.3
Attn. Tune 80.8 71.7 87.2 87.8 81.9
Model Parameter Averaging (MPA) 78.7 73.4 88.4 87.6 82.0
CLIPOOD (Shu et al., 2023) 87.9 77.8 94.1 93.6 88.3
MIRO + SWAD 85.6 79.9 93.0 92.1 87.6
MIRO + MPA 85.6 80.5 93.0 92.3 87.9
Table 8: Per-domain breakdown for TerraIncognita (DomainBed-OOP)
Method L100 L38 L43 L46 Avg
CORAL 13.9 8.1 21.1 21.1 16.0
SAGM 17.2 7.8 20.2 32.2 19.3
ERM* 13.1 9.6 22.9 20.5 16.5
LP-FT 16.7 5.1 25.6 44.9 23.1
SWAD 20.7 8.2 26.9 28.2 21.0
MIRO 19.4 6.6 30.0 17.9 18.5
VL2V-SD 9.3 5.6 20.2 28.4 15.9
Attn. Tune 20.5 10.7 24.8 25.3 20.3
Model Parameter Averaging (MPA) 21.7 8.6 30.9 27.6 22.2
CLIPOOD (Shu et al., 2023) 37.9 11.3 21.6 8.9 19.9
MIRO + SWAD 22.5 5.9 30.3 25.9 21.1
MIRO + MPA 24.8 9.6 32.5 32.8 24.9
Table 9: Per-domain breakdown for TerraIncognita (DomainBed-IP)
Method L100 L38 L43 L46 Avg
CORAL 46.4 35.3 56.9 32.8 42.9
SAGM 55.1 38.7 52.5 29.5 44.0
ERM* 37.4 38.0 55.2 37.5 42.0
LP-FT 50.7 36.5 58.1 18.2 40.9
SWAD 60.1 41.7 65.7 43.5 52.7
MIRO 69.5 50.2 69.7 46.1 58.9
VL2V-SD 52.8 38.4 57.6 43.8 48.1
Attn. Tune 52.4 44.0 68.0 47.5 53.0
Model Parameter Averaging (MPA) 63.8 41.4 67.9 44.7 54.4
CLIPOOD (Shu et al., 2023) 77.7 56.4 69.0 51.0 63.5
MIRO + SWAD 68.9 54.8 73.4 51.2 62.1
MIRO + MPA 69.4 55.9 73.6 51.4 62.6
Table 10: Per-domain breakdown for TerraIncognita (DomainBed-All)
Method L100 L38 L43 L46 Avg
CORAL 43.6 32.7 50.4 31.8 39.6
SAGM 51.9 35.8 46.6 29.7 41.0
ERM* 35.3 35.3 49.3 36.0 39.0
LP-FT 47.8 33.5 52.1 20.5 38.5
SWAD 56.7 38.5 58.6 42.2 49.0
MIRO 65.2 46.0 62.4 43.7 54.3
VL2V-SD 49.1 35.3 50.8 42.5 44.4
Attn. Tune 49.7 40.8 60.1 45.6 49.1
Model Parameter Averaging (MPA) 60.1 38.3 61.1 43.2 50.7
CLIPOOD (Shu et al., 2023) 74.3 52.1 60.4 47.4 58.5
MIRO + SWAD 64.9 50.1 65.5 49.0 57.4
MIRO + MPA 65.6 51.5 66.0 49.8 58.2
Table 11: Comparison of Methods for DomainNet Dataset (DomainBed-OOP)
Method Clp Inf Pnt Qkdr Real Skt Avg
CORAL 34.9 19.2 26.9 5.7 24.6 22.8 22.3
SAGM 36.4 19.4 25.1 6.1 26.8 24.2 23.0
ERM* 36.2 19.0 23.4 5.4 26.6 23.4 22.3
LP-FT 34.7 20.9 25.2 6.1 25.9 23.2 22.7
SWAD 42.5 28.3 33.4 7.6 30.2 29.5 28.6
MIRO 42.5 37.0 27.9 3.5 33.5 25.8 28.4
VL2V-SD 43.0 43.2 34.8 3.0 37.1 29.5 31.8
Attn. Tune 40.4 29.1 29.2 4.3 31.3 26.6 26.8
Model Parameter Averaging (MPA) 42.3 30.9 34.9 7.4 31.6 30.7 29.6
CLIPOOD (Shu et al., 2023) 39.2 50.2 39.5 2.1 40.9 31.3 33.9
MIRO + SWAD 46.1 40.4 34.0 4.4 35.2 31.7 32.0
MIRO + MPA 46.2 42.6 36.0 4.4 36.0 33.3 33.1
Table 12: Comparison of Methods for DomainNet Dataset (DomainBed-IP)
Method Clp Inf Pnt Qkdr Real Skt Avg
CORAL 80.3 52.3 71.1 33.2 71.7 71.2 63.3
SAGM 81.7 52.8 70.2 34.5 73.4 73.1 64.3
ERM* 80.6 53.4 69.0 31.3 73.0 71.6 63.1
LP-FT 80.0 56.5 70.5 34.0 73.9 71.5 64.4
SWAD 85.5 67.1 80.2 41.3 80.1 79.8 72.3
MIRO 85.4 77.7 79.1 34.6 83.2 74.2 72.4
VL2V-SD 88.3 85.6 84.6 38.0 88.3 84.0 78.1
Attn. Tune 83.5 66.1 77.3 32.6 80.3 75.5 69.2
Model Parameter Averaging (MPA) 85.8 69.7 81.6 41.4 81.4 81.4 73.6
CLIPOOD (Shu et al., 2023) 86.6 85.4 87.6 38.0 91.0 84.7 78.9
MIRO + SWAD 88.2 81.7 83.9 40.8 85.8 81.9 77.0
MIRO + MPA 88.1 83.7 85.5 41.3 86.5 84.0 78.2
Table 13: Comparison of Methods for DomainNet Dataset (DomainBed-All)
Method Clp Inf Pnt Qkdr Real Skt Avg
CORAL 72.7 27.1 58.0 20.1 67.1 58.8 50.6
SAGM 74.1 27.4 56.9 20.9 68.9 60.6 51.5
ERM* 73.1 27.4 55.7 19.0 68.5 59.3 50.5
LP-FT 72.4 29.3 57.1 20.7 69.2 59.2 51.3
SWAD 78.0 36.0 66.1 25.2 75.2 66.7 57.9
MIRO 78.0 42.9 64.0 20.0 78.3 61.7 57.5
VL2V-SD 80.5 47.8 69.5 21.6 83.1 69.8 62.0
Attn. Tune 76.1 36.0 63.0 19.2 75.4 62.9 55.4
Model Parameter Averaging (MPA) 78.3 37.9 67.4 25.2 76.5 68.2 58.9
CLIPOOD (Shu et al., 2023) 78.4 52.9 72.6 21.2 85.9 70.7 63.6
MIRO + SWAD 80.8 45.6 68.8 23.7 80.7 68.7 61.4
MIRO + MPA 80.7 47.2 70.4 23.9 81.4 70.6 62.4
Table 14: Per-domain breakdown for VLCS (DomainBed-OOP)
Method Caltech101 LabelMe SUN09 VOC2007 Avg
CORAL 93.3 54.6 72.9 75.3 74.0
SAGM 80.0 58.8 76.5 78.0 73.3
ERM* 100.0 57.1 73.3 75.3 76.4
LP-FT 80.0 55.4 71.8 75.7 70.7
SWAD 100.0 51.2 79.1 77.7 77.0
MIRO 86.7 52.6 82.9 72.7 73.7
VL2V-SD 100.0 55.0 80.4 81.1 79.1
Attn. Tune 86.7 61.0 77.9 79.0 76.1
Model Parameter Averaging (MPA) 100.0 53.6 83.6 80.9 79.5
CLIPOOD (Shu et al., 2023) 100.0 55.0 84.9 82.8 80.7
MIRO + SWAD 100.0 56.7 84.7 79.9 80.3
MIRO + MPA 100.0 54.6 83.4 77.4 78.9
Table 15: Per-domain breakdown for VLCS (DomainBed-IP)
Method Caltech101 LabelMe SUN09 VOC2007 Avg
CORAL 98.2 83.5 84.5 79.8 86.5
SAGM 95.9 84.3 86.3 85.5 88.0
ERM* 96.9 83.7 86.1 74.6 85.3
LP-FT 97.1 83.2 82.8 80.7 86.0
SWAD 97.9 82.9 88.4 84.8 88.5
MIRO 98.9 83.3 93.5 88.2 91.0
VL2V-SD 99.5 83.9 91.8 94.6 92.4
Attn. Tune 98.0 84.7 85.7 86.5 88.7
Model Parameter Averaging (MPA) 98.4 83.0 92.6 88.8 90.7
CLIPOOD (Shu et al., 2023) 98.9 83.6 93.5 94.0 92.5
MIRO + SWAD 97.6 83.4 92.6 90.6 91.1
MIRO + MPA 98.0 83.2 92.9 90.1 91.0
Table 16: Per-domain breakdown for VLCS (DomainBed-All)
Method Caltech101 LabelMe SUN09 VOC2007 Avg
CORAL 98.2 66.2 72.7 77.0 78.5
SAGM 95.8 68.8 76.2 80.8 80.4
ERM* 96.9 67.8 73.0 73.9 77.9
LP-FT 97.0 66.4 71.4 77.4 78.0
SWAD 98.0 64.5 77.7 80.3 80.1
MIRO 98.7 65.3 81.2 79.3 81.1
VL2V-SD 99.5 66.6 78.8 86.0 82.7
Attn. Tune 97.9 70.3 77.2 81.9 81.8
Model Parameter Averaging (MPA) 98.4 65.7 81.3 83.7 82.3
CLIPOOD (Shu et al., 2023) 98.9 66.7 81.6 86.6 83.4
MIRO + SWAD 97.6 66.4 81.6 82.6 82.0
MIRO + MPA 98.0 67.4 82.2 83.6 82.8
Table 17: Per-domain breakdown for PACS (DomainBed-OOP)
Method Art Cartoon Photo Sketch Avg
CORAL 80.4 84.3 89.5 42.2 74.1
SAGM 74.4 78.1 89.5 54.7 74.2
ERM* 75.9 86.8 91.6 53.2 76.9
LP-FT 74.2 85.9 97.9 56.3 78.6
SWAD 84.6 81.6 95.8 54.3 79.1
MIRO 94.2 94.1 97.9 52.8 84.7
VL2V-SD 93.5 96.5 97.9 52.2 85.0
Attn. Tune 94.0 91.4 95.8 55.7 84.2
Model Parameter Averaging (MPA) 94.0 88.9 97.9 50.2 82.7
CLIPOOD (Shu et al., 2023) 96.0 96.8 97.9 58.3 87.2
MIRO + SWAD 96.0 94.9 100.0 50.7 85.4
MIRO + MPA 95.3 95.4 100.0 60.3 87.8
Table 18: Per-domain breakdown for PACS (DomainBed-IP)
Method Art Cartoon Photo Sketch Avg
CORAL 83.3 89.8 91.8 72.1 84.3
SAGM 89.7 93.0 94.7 82.9 90.1
ERM* 88.7 93.0 91.5 75.2 87.1
LP-FT 85.3 93.7 97.9 84.2 90.3
SWAD 96.4 93.0 97.8 91.0 94.6
MIRO 99.2 96.7 99.9 94.4 97.6
VL2V-SD 99.6 98.9 99.9 93.6 98.0
Attn. Tune 97.3 97.6 98.9 91.7 96.4
Model Parameter Averaging (MPA) 98.6 95.8 98.2 89.2 95.4
CLIPOOD (Shu et al., 2023) 99.4 99.3 99.9 92.0 97.7
MIRO + SWAD 99.1 97.9 99.9 93.3 97.6
MIRO + MPA 99.0 98.5 100.0 94.7 98.1
Table 19: Per-domain breakdown for PACS (DomainBed-All)
Method Art Cartoon Photo Sketch Avg
CORAL 81.4 89.2 91.7 70.5 83.2
SAGM 83.4 90.4 94.6 81.4 87.5
ERM* 83.0 92.1 91.5 74.0 85.2
LP-FT 80.5 92.7 97.9 82.7 88.4
SWAD 91.5 91.2 97.8 89.0 92.4
MIRO 97.1 96.4 99.8 92.2 96.4
VL2V-SD 97.7 98.5 99.9 91.3 96.9
Attn. Tune 96.4 96.5 98.8 89.7 95.4
Model Parameter Averaging (MPA) 97.2 94.8 98.2 87.1 94.3
CLIPOOD (Shu et al., 2023) 98.2 98.8 99.8 90.2 96.8
MIRO + SWAD 98.0 97.4 99.9 91.0 96.6
MIRO + MPA 97.8 98.0 99.9 92.9 97.2
Figure 14: Class distribution of DomainNet-IP. Zoom in on pdf for best viewing.
Figure 15: Class distribution of DomainNet-OOP. Zoom in on pdf for best viewing.
Figure 16: Class distribution of DomainNet before splitting. Zoom in on PDF before viewing.
Figure 17: Class distribution of OfficeHome-IP. Zoom in on pdf for best viewing.
Figure 18: Class distribution of OfficeHome-OOP. Zoom in on pdf for best viewing.
Figure 19: Class distribution of OfficeHome before Splitting. Zoom in on pdf for best viewing.
Figure 20: Class distribution of PACS-IP.
Figure 21: Class distribution of PACS-OOP.
Figure 22: Class distribution of PACS before splitting.
Figure 23: Class distribution of TerraIncognita-IP
Figure 24: Class distribution of TerraIncognita-OOP
Figure 25: Class distribution of TerraIncognita before splitting.
Figure 26: Class distribution of VLCS-IP
Figure 27: Class distribution of VLCS-OOP
Figure 28: Class distribution of VLCS before splitting.